OpenStudy (anonymous):

for what value of k does the linear system have no solution? kx+3y=11 3x-y=11

OpenStudy (anonymous):

For no solution: \[ \frac k 3 = -\frac 3 1 \neq \frac{11}{11} \]

OpenStudy (anonymous):

So, \(k =-9 \)

OpenStudy (anonymous):

oh well thank that what i needed!!!!!

OpenStudy (anonymous):

Glad to help.

OpenStudy (anonymous):

for what value of k does the linear system have no solution? 2x+ky=11 3x-y=11